OpenStudy (anonymous):

The United States Department of Agriculture reports that 15% Americans are now on food stamps. If there are about 300 million Americans, how many millions of people are on food stamps?

OpenStudy (mathstudent55):

15% of 300,000,000 = 15 * 1/100 * 300,000,000 = 0.15 * 300,000,000

OpenStudy (anonymous):

so its 300,000,000?

OpenStudy (mathstudent55):

What is 0.15 times 300,000,000?

OpenStudy (anonymous):

oh lol 45,000,000

OpenStudy (anonymous):

Thank you!!!

OpenStudy (mathstudent55):

Right, but the question is how many millions ..., so answer is 45, not 45 million.

OpenStudy (anonymous):

alrighty then. Thank you once again!

OpenStudy (mathstudent55):

wlcm